### v1

# Evaluating Scoring Bias in LLM-as-a-Judge

Abstract
The remarkable performance of Large Language Models (LLMs) gives rise to“LLM-as-a-Judge”, where LLMs are employed as evaluators for complex tasks. Moreover, it has been widely adopted across fields such as Natural Language Processing (NLP), preference learning, and various specific domains. However, there are various biases within LLM-as-a-Judge, which adversely affect the fairness and reliability of judgments. Current research on evaluating or mitigating bias in LLM-as-a-Judge predominantly focuses on comparison-based evaluations, while systematic investigations into bias in scoring-based evaluations remain limited. Therefore, we define scoring bias in LLM-as-a-Judge as the scores differ when scoring judge models are bias-related perturbed, and provide a well-designed framework to comprehensively evaluate scoring bias. We augment existing LLM-as-a-Judge benchmarks through data synthesis to construct our evaluation dataset and design multi-faceted evaluation metrics. Our experimental results demonstrate that the scoring stability of existing judge models is disrupted by scoring biases. Further exploratory experiments and discussions provide valuable insights into the design of scoring prompt templates and the mitigation of scoring biases on aspects such as score rubrics, score IDs, and reference answer selection. [^1] In the interest of protecting our intellectual property, specific technical details have been redacted from this preliminary version. The complete and unredacted manuscript is forthcoming.
Introduction
Large Language Models (LLMs), with outstanding performance, have been widely applied to Natural Language Generation (NLG) tasks such as summarization (Zhang et al. 2024) question answering (Robinson, Rytting, and Wingate 2022), and domain-specific dialogues (Wang et al. 2024a; Singhal et al. 2025). In these tasks, the model’s responses require find-grained evaluation of various aspects such as quality, factuality, and instruction-following, thereby providing valuable insights for model improvement. Traditional statistical NLG evaluation metrics, such as BLEU (Papineni et al. 2002) and ROUGE (Lin 2004), evaluate the generated text by n-gram overlap between the text and the golden answer. The above methods exhibit two primary limitations: (1) They focus solely on the similarity between the generated text and the golden answer, not quantifying intrinsic attributes of the text such as writing style, factuality, and educational value. (2) The resulting metrics lack interpretability, making it difficult to provide guidance for the model upgrade. There is an urgent need for fine-grained, multifaceted, and interpretable evaluation metrics and methods tailored to NLG tasks and LLM applications. With remarkable capabilities in instruction understanding and response generation, LLMs can provide diverse forms of evaluation output, such as scoring, comparison, and ranking. In addition, they can conduct evaluations along specified dimensions and generate critics for judgments, thus enhancing the interpretability of the evaluation. This line of research, referred to as “LLM-as-a-Judge” (Gu et al. 2024; Li et al. 2024a), has recently emerged as a prominent focus.
Figure 1: Our evaluation framework to measure the scoring bias in LLM-as-a-Judge.
Robustness is a critical factor in ensuring the fairness and reliability of LLM judgments, which requires that, under identical judge rubrics, LLM judges should provide consistent judgments for the same evaluation target. However, recent work (Park et al. 2024) indicates that LLM judges are subject to various biases, which undermine the self-consistency of their judgments. Many studies have sought to identify the types of bias that shift in judgments. (Wang et al. 2024b) finds that the judgments of the candidate responses can be easily hacked by altering their position in the context. (Wataoka, Takahashi, and Ri 2024) measures the self-preference bias in LLM-as-a-Judge, which refers to the tendency of LLM judges to overestimate the quality of their own outputs. (Wei et al. 2024) considers length bias as LLM judges favor longer responses or shorter responses.
Although existing studies systematically investigated various biases in LLM-as-a-Judge, several pressing issues remain to be addressed. Firstly, while some work (Wang et al. 2024b; Ye et al. 2024a) has addressed biases such as positional bias and refinement awareness in scoring-based LLM judges, many existing research remains focused on biases in comparative LLM judges. Existing work (Ye et al. 2024a) attempts to investigate specific biases in scoring tasks, such as self-enhancement bias and refinement-aware bias. However, there is still a lack of comprehensive research on biases in scoring-based LLM-as-a-Judge. Secondly, the biases defined in existing work are predominantly related to the attributes or presentation forms of the evaluation targets, such as position, length, and response model. However, variation inherent in the prompt of LLM-based judge itself may also impact the self-consistency of judgments, which lacks a systematic investigation in current research.
To solve the above problems, we propose scoring bias, defined as the scoring shifts of LLM-as-a-Judge under perturbations of the scoring prompt. In this work, we propose three types of scoring bias: score rubrics order bias, score IDs bias, and reference answer score bias, and design a framework to systematically evaluate those biases. We select four existing LLM-as-a-Judge benchmarks as evaluation datasets and develop a series of metrics to quantify the robustness of LLM judges against the score bias. Our key contributions can be summarized as follows:
  • We pioneer a systematic evaluation of bias in scoring-based LLM-as-a-Judge with definition of three types of scoring bias and five perturbation methods.
  • Through automatic data synthesis, we expand existing LLM-as-a-Judge benchmarks into our evaluation dataset without human labor costs. Furthermore, we evaluate scoring bias from multiple perspectives through a comprehensive design of evaluation metrics.
  • Our comprehensive experiments and in-depth analysis find that scoring bias affects the robustness and reliability of various existing judge models. We also provide valuable recommendations for the design of scoring prompts and the mitigation of scoring bias.
Related Work
LLM-as-a-Judge
(Gu et al. 2024) defines LLM-as-a-Judge as the task in which an LLM generates an evaluation result in any form given input data and context. Common forms of LLM judgment include score generation (Gao et al. 2023), pairwise comparison (Qin et al. 2024), ranking (Li, Patel, and Du 2023), multiple-choice selection (Li et al. 2024b), and critic generation (Ke et al. 2024). Improvement strategies for LLM-as-a-Judge include meticulous design of evaluation prompts (Liu et al. 2023), improvement of LLM’s abilities such as fine-tuning (Wang et al. 2024c), multi-LLM collaboration pipelines such as feedback mechanism (Xu et al. 2023), and output optimization strategies such as multi-output ensemble (Sottana et al. 2023). The performance of LLM-as-a-Judge is usually measured using metrics that represent its alignment with human annotators or advanced LLM judgments such as agreement (Thakur et al. 2024), Cohen’s Kappa, and Spearman’s correlation (Bai et al. 2023). LLM-as-a-Judge is widely applied to evaluation (Lin and Chen 2023), alignment (Bai et al. 2022), retrieval (Sun et al. 2023), reasoning (Gao et al. 2024a), and other specific domains (Ryu et al. 2023; Babaei and Giudici 2024). LLM-as-a-Judge faces challenges in bias and vulnerability, which significantly compromise the fairness and reliability of LLM judges. (Li et al. 2024a)
Bias in LLM-as-a-Judge
(Ye et al. 2024a) defines the prompt for LLM judges as three components: system instruction, question, and response. Based on this, they define bias in LLM-as-a-Judge as that the judgments differ when system instruction or responses is bias-related modified. Common types of bias in LLM-as-a-Judge include position bias, length bias, self-enhancement bias, etc. Those biases are often incorporated into the evaluation scope of existing LLM-as-a-Judge benchmarks. (Zheng et al. 2023; Wei et al. 2024) Currently, the types of bias still remain a focal point of investigation, and numerous benchmarks are proposed to evaluate bias in LLM-as-a-Judge. OffsetBias (Park et al. 2024) points out six types of bias in LLM-as-a-Judge including length bias, concreteness bias, empty reference bias, and so on. CALM (Ye et al. 2024a) quantifies and analyzes 12 types of bias in LLM judges. However, the biases defined in current work are predominantly focused on comparative evaluation tasks, with a lack of systematic research and definitions specifically addressing biases in scoring-based evaluation tasks. Similar to research on adversarial attacks, the evaluation metrics for bias in LLM-as-a-Judge are often based on the consistency between bias-related modified judgments and the original judgments, or on alignment with golden answers. Some studies focus on mitigating bias in LLM-as-a-Judge and aim to develop more robust LLM judges. (Liu et al. 2024c) introduces Pairwise-Preference Search (PAIRS) to mitigate biases in LLM evaluators. (Wang et al. 2024b) proposes three methods to calibrate positional bias in LLM-as-a-Judge, including multiple evidence calibration, balanced position calibration, and human-in-the-loop calibration. Our evaluation results on scoring bias will serve as a reference to guide future research in mitigating scoring biases in judge models and enhancing the robustness of scoring-based evaluations.
Evaluation Method
Figure 2: The scoring prompt template with four components from (Kim et al. 2024c). The red text represents the components may be perturbed.
Figure 1 shows our evaluation framework, inspired by (Ye et al. 2024a).
Dataset Data Size Human Judgment Reference Answer
BiGGen Bench 2,780
FLASK 2,001
MT Bench 320
Vicuna Bench 320
Table 1: Statics of our evaluation datasets.
We evaluate the scoring bias in LLM-as-a-Judge by perturbing the scoring prompt and observing whether the scoring judgments are affected, which follows (Kim et al. 2024c).
Bias Types
We define three types of bias targeting the key components related to scoring in LLM judges:
  • Score Rubrics Order
  • Score IDs
  • Reference Answer Score
Datasets
Following previous work (Kim et al. 2024c; Chiang, Lee, and Lukasik 2025), we select four LLM-as-a-Judge datasets for scoring bias evaluation:
  • BiGGen Bench (Kim et al. 2024b): This is a principled generation benchmark with instance-specific evaluation criteria. We use its test set for non-multilingual capabilities with 2,780 responses from 695 instances that are evaluated by human annotators.
  • FLASK (Ye et al. 2024b): This is a fine-grained evaluation benchmark with 200 test prompts, 12 score rubrics, and 2,001 responses. It also includes scores labeled by human annotators.
  • MT Bench (Zheng et al. 2023): This is a multi-turn chat benchmark with 80 test prompts. (Kim et al. 2024a) write 80 hand-crafted score rubrics and generate 320 responses from WizardLM-13B (Xu et al. 2024), Vicuna-13B (Chiang et al. 2023), Llama-2-Chat-13B (Touvron et al. 2023), and GPT-3.5-Turbo-0613 (OpenAI 2023a).
  • Vicuna Bench (Chiang et al. 2023): This is a single-turn chat benchmark with 80 test prompts.(Kim et al. 2024a) also write 80 hand-crafted score rubrics and generate 320 responses from WizardLM-13B, Vicuna-13B, Llama-2-Chat-13B, and GPT-3.5-Turbo-0613.
Model BiGGen Bench FLASK
$\rho$ $r$ $\rho$ $r$
GPT-4.1 0.6048 0.6360 0.6401 0.6823
GPT-4o 0.6047 0.6340 0.6340 0.6557
GPT-4 0.5741 0.6045 0.6207 0.6862
Table 2: Correlations between LLM judgement pipeline and human when generating golden scores. $\rho$ is Spearman’s rank correlation coefficient. $r$ is Pearson’s correlation coefficient. The GPT-4.1 and GPT-4o scores are generated by our work, the GPT-4 scores are derived from the original datasets.
The detailed statics of our selected datasets are shown in Table 1. For advanced LLM scores, the above datasets have evaluation results obtained from GPT-4 (OpenAI 2023b). Due to the earlier release of GPT-4, we use more recent models GPT-4.1 (OpenAI 2025) and GPT-4o (OpenAI 2024) to obtain golden scores. Inspired by (Kim et al. 2024a, c), we repeat the scoring three times through a judge prompt with a reference answer that marked a score of 5 and make a majority vote to obtain the final score from LLM evaluators. Table 2 shows the correlation between advanced LLM scores and human scores. The results show that the GPT-4.1 scores exhibit the highest correlation with human scores, so we select the GPT-4.1 scores as the golden LLM scores.
Each instance in BiGGen Bench and FLASK is accompanied by a reference answer with score 5. To measure the impact of the reference answer score on the judgment robustness, it is necessary to additionally generate reference answers with scores ranging from 1 to 4. We propose a generation-review pipeline to synthesize reference answers for each sample in FLASK, as shown in Figure 3. Given an instruction and score rubrics, the response model generates the response according to the description from score 1 to 4. Then a review model checks whether the response matches the corresponding score description. If match, we accept the generated responses; otherwise, we rerun the pipeline until a match is achieved. We alternately apply GPT-4.1 and GPT-4o as response model and review model in the pipeline.
Metrics
Figure 3: The proposed generation-review pipeline for synthesizing reference answers.
To evaluate the scoring bias in LLM-as-a-Judge, we consider the impact or perturbation methods on scoring accuracy and scoring tendency.
Experiments
Model Bias Perturbation BiGGen Bench FLASK MT Bench Vicuna Bench
Human GPT-4.1 Human GPT-4.1 GPT-4.1 GPT-4.1
$\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$ $\rho$ $r$
GPT-4o - - 0.5685 0.6165 0.7572 0.8059 0.5529 0.5914 0.7369 0.7676 0.7578 0.7724 0.6862 0.7696
Rubrics Order Descending 0.5852 0.6245 0.7787 0.8233 0.5754 0.6080 0.7602 0.7760 0.7479 0.7597 0.6734 0.7635
Random 0.5963 0.6313 0.7698 0.8124 0.5592 0.5893 0.7420 0.7573 0.7673 0.7838 0.6701 0.7504
Score IDs Letter 0.5641 0.6034 0.7595 0.7994 0.5558 0.5977 0.7341 0.7609 0.7317 0.7454 0.6539 0.7712
Roman numerals 0.5754 0.6171 0.7709 0.8148 0.5535 0.5912 0.7380 0.7641 0.7725 0.7786 0.7073 0.7821
DeepSeek-V3 - - 0.5314 0.5748 0.6775 0.7394 0.5163 0.5660 0.6711 0.7063 0.6323 0.6593 0.4313 0.5443
Rubrics Order Descending 0.5356 0.5812 0.6611 0.7211 0.5181 0.5554 0.6555 0.6858 0.6183 0.6436 0.4993 0.5854
Random 0.5136 0.5609 0.6547 0.7125 0.4999 0.5470 0.6521 0.6752 0.6178 0.6460 0.5086 0.5482
Score IDs Letter 0.5378 0.5689 0.6803 0.7179 0.5314 0.5755 0.6663 0.6999 0.6489 0.6715 0.4832 0.5657
Roman numerals 0.5164 0.5636 0.6688 0.7195 0.5066 0.5542 0.6750 0.7109 0.6558 0.6642 0.4416 0.5285
Qwen3-32B - - 0.5421 0.5855 0.7328 0.7715 0.5323 0.5695 0.6549 0.6807 0.6959 0.7053 0.6419 0.7188
Rubrics Order Descending 0.5572 0.5989 0.7443 0.7862 0.5345 0.5657 0.6586 0.6839 0.6843 0.6989 0.6205 0.7020
Random 0.5528 0.5938 0.7404 0.7824 0.5115 0.5456 0.6537 0.6748 0.6554 0.6688 0.6416 0.7252
Score IDs Letter 0.5353 0.5851 0.7280 0.7770 0.5393 0.5724 0.6650 0.6975 0.6795 0.6930 0.6157 0.6808
Roman numerals 0.5139 0.5629 0.6992 0.7412 0.5257 0.5781 0.6630 0.6980 0.6360 0.6418 0.6072 0.7126
Qwen3-8B - - 0.5092 0.5308 0.6522 0.6492 0.5035 0.5309 0.6464 0.6616 0.6238 0.6342 0.6639 0.7284
Rubrics Order Descending 0.5241 0.5456 0.6458 0.6559 0.4885 0.5339 0.6315 0.6566 0.6005 0.6135 0.5500 0.6647
Random 0.4869 0.5025 0.6091 0.5950 0.4778 0.5004 0.6061 0.6216 0.5949 0.5983 0.5135 0.6053
Score IDs Letter 0.4853 0.5117 0.6410 0.6665 0.4873 0.5244 0.6334 0.6586 0.5985 0.6126 0.5812 0.6890
Roman numerals 0.4801 0.5120 0.6231 0.6445 0.4606 0.5087 0.6177 0.6468 0.6016 0.5975 0.6111 0.7178
Mistral-24B - - 0.5059 0.5495 0.6910 0.7201 0.4950 0.5385 0.6303 0.6593 0.5943 0.5762 0.6116 0.6917
Rubrics Order Descending 0.5369 0.5761 0.6987 0.7213 0.4922 0.5327 0.6362 0.6534 0.6104 0.6102 0.6137 0.7082
Random 0.5053 0.5450 0.6569 0.6841 0.4756 0.5042 0.6212 0.6338 0.5603 0.5602 0.5461 0.6418
Score IDs Letter 0.5114 0.5632 0.6685 0.7084 0.4880 0.5339 0.6311 0.6594 0.5375 0.5485 0.5558 0.6608
Roman numerals 0.5102 0.5590 0.6640 0.6962 0.4957 0.5494 0.6533 0.6872 0.5863 0.5907 0.5903 0.6905
Table 3: Correlation between golden scores and judge model scores with bias in score rubrics order and score IDs. For each judge model, the first row is the metrics without bias, the remaining rows is the metrics under different types of bias. The differences between results with and without bias reflect the extent to which bias impacts the scoring robustness. DeepSeek-V3 is DeepSeek-V3-671B. Mistral-24B is Mistral-Small-24B-Instruct. $\rho$ is Spearman’s rank correlation coefficient. $r$ is Pearson’s correlation coefficient.
Setup
Judge Models.
Previous work (Gao et al. 2024b; Li et al. 2024c; Liu et al. 2024b) has reached a consensus that judge models should be stronger models to ensure evaluation reliability. (Ye et al. 2024a) However, as an exploratory study, we think that capability and size of the judge model should also be considered as influence factors. Therefore, we select diverse judge models that include closed-source, open-source, large-size, and small-size models. They are GPT-4o, DeepSeek-V3-671B (Liu et al. 2024a), Qwen3-32B, Qwen3-8B (Yang et al. 2025), and Mistral-Small-24B-Instruct-2501 (MistralAI 2025). Those judge models we selected differ from the models used to generate response in the evaluation datasets, thus mitigating potential self-enhancement bias (Zheng et al. 2023; Ye et al. 2024a).
Judge Prompt.
The prompt we use for scoring-based judgment is shown in Figure 2, following (Kim et al. 2024c). The unperturbed prompt is the one with ascending score rubrics, Arabic numeral score IDs, and no reference answer.
Hyperparameters.
As our goals for scoring bias evaluations, we aim to ensure the stability of the model’s generative capabilities. Therefore, we set the judge models’ temperature to 0 while scoring.
Main Results
Judge models are sensitive to scoring biases in input prompt templates.
Table 3 and Figure 4 show the correlation of judge model scores and golden scores on all perturbation methods. All judge models we evaluated, even for an advanced model such as GPT-4, fluctuations in score results are observed under different biases, indicating that scoring bias has a significant impact on the robustness of LLM-as-a-Judge. Furthermore, the impact exhibits certain differences depending on the choice of judge model:
(1) The capability and size of the judge models are critical factors influencing the robustness against scoring bias. Consistent with existing consensus (Gao et al. 2024b; Li et al. 2024c; Liu et al. 2024b), our experimental results indicate that judge models with advanced capability and larger size tend to exhibit higher reliability on scoring. As shown in Table 3, for example, across all perturbation methods and datasets, the fluctuation in correlation coefficients between GPT-4o and the golden scores remains with in 0.03, except in introduction of a reference answer scored 5. By comparing the results of Qwen3-32B and Qwen3-8B, it is evident that judge models with larger parameter sizes demonstrate greater robustness in scoring. In comparison with the GPT-4.1 scores in BiGGen-Bench and Vicuna Bench, the correlation coefficients of the Qwen3-8B scores exhibit maximum fluctuations exceeding 0.05 under score rubrics order bias and score IDs bias. With reference answer score bias, the maximum fluctuation reaches nearly 0.2. However, the changes in the correlation coefficient of Qwen3-32B under biases are generally within 0.03. Therefore, more capable judge models should be prioritized in scoring evaluations.
Figure 4: Spearman’s correlation between human annotator scores and judge model scores with bias in reference answer score on FLASK.
(2) Different judge models exhibit varying sensitivity to scoring bias. The impact of bias on the scoring accuracy of judge models can be either positive or negative. However, this impact is not fixed; different judge models have different trends in scoring accuracy when affected by the same bias. As Table 3 shows, for example, in most cases, the correlation between GPT-4o scores and golden scores increases when Roman numerals are used as scoring IDs. But when DeepSeek-V3-671B, Qwen3-32B and Qwen3-8B are employed as judge models, the use of Roman numerals generally exerts a negative effect on their scoring accuracy. An interesting phenomenon is that the impact of reference answer score bias on the judge models we evaluated appears to be consistent, which will be further analyzed in subsequent sections. Our experimental results can serve as a valuable reference for the selection of judge models and the design of scoring prompt templates.
(3) Judge models have their own scoring tendencies. Figure 5 presents the scoring tendencies of different judge models under various scoring biases on the same dataset FLASK in the form of heatmaps. It is obvious that each judge model exhibits its own specific scoring preferences. Under unbiased conditions, DeepSeek-V3-671B assigns a score of 5 to more than half of the instruction-response pairs. Compared to the human scoring distribution, Mistral-Small-24B-Instruct has a stronger tendency to score a 4 to the samples. Other judge models show smoother scoring distribution shifts, but certain scores are still preferred. GPT-4o demonstrates a preference for score 4, while Qwen3-32B favors score 5. An examination of Figure 5 reveals that score rubrics order bias and score IDs bias have a certain impact on the scoring tendencies of the judge models; however, they do not fundamentally alter the overall score distribution. An exception is Qwen3-8B, whose score distribution is more significantly influenced by biases, possibly due to its small size. It is particularly evident that the reference answer score bias exerts a substantial impact on the score distribution, which will be analyzed in detail in the following sections. Our experimental results show that judge models inherently exhibit biases in their scoring tendencies. Therefore, debiasing the judge models is crucial to ensuring the fairness and reliability of the evaluation.
Figure 5: Heatmaps of judge models’ score distribution with all bias types on FLASK. The horizontal axis is the score, where -1 represents the failed judgment. The vertical axis is the perturbation method. The top lines of all heatmaps are the human scoring tendencies for comparison, and the second lines are results from unbiased judge models.
Exploratory Experiments and Discussion
For advanced or large size judge models, application of scoring prompt that differ from usual preferences are worth to be considered.
In almost all scoring-based evaluation tasks, Arabic numerals are commonly used as scoring IDs. The score rubrics order typically follows a sequential order from score 1 to score 5. These choices align with human writing conventions and cognitive preferences in the design of scoring prompts. However, our experimental results shown in Table 3 indicate that the scoring prompts designed to align with the usual preferences do not necessarily produce the most accurate scoring results. As mentioned earlier, the performance of the GPT-4o judge model in scoring evaluations improves when Roman numerals are used as scoring IDs. In more than half of the cases, using letter grades as scoring Ids improves the scoring accuracy of DeepSeek-V3-671B. Regarding the score rubrics order bias, adopting a descending order improves the scoring accuracy of GPT-4o, Qwen3-32B, and Mistral-Small-24B-Instruct in most instances. Intuitively, employing a random scoring rubrics order often tends to negatively impact the scoring accuracy of all judge models. Based on the above analysis, we recommend incorporating certain intentional deviations from conventional practices in the design of scoring prompts.
The reference answer score bias perform intriguing impacts on the robustness of scoring-based evaluation and play a critical role.
The impact of reference answer score bias on the scoring accuracy and scoring tendency of the judge model is presented in Figure 4 and Figure 5. The experimental results indicate that, compared to the score rubrics order bias and the score IDs bias, the influence of the reference answer score bias on scoring robustness is both more pronounced and demonstrates stronger regularity. Figure 4 illustrates the impact of introducing reference answers with different scores in the input prompt on the scoring accuracy of the judge model. For GPT-4o, DeepSeek-V3-671B, and Mistral-Small-24B-Instruct, the scoring accuracy in most cases improves when reference answers are introduced. However, for the two Qwen judge models, only reference answers with scores of 4 or 5 contribute to enhancing scoring performance. Figure 4 also reveals an interesting phenomenon: the higher the score of the reference answer, the better the scoring accuracy of the judge model. Although DeepSeek-V3-671B, Qwen3-32B, and Mistral-Small-24B-Instruct do not strictly adhere to this pattern, the correlation between their scores and human scores similarly exhibits an upward trend as the reference answer score increases.
Figure 5 shows that the score of the reference answer has a significant impact on the scoring tendency of the judge model. Specifically, the scoring results from DeepSeek-V3-671B, Qwen3-32B, and Qwen3-8B appear to be highly influenced by the score of the reference answer. When the score of the reference answer falls within the range of 1 to 4, these models tend to assign similar scores to the evaluated instruction-response pairs. Interestingly, when a reference answer with a score of 5 is introduced, the scoring behavior of the judge model appears to become more rational. This is particularly notable for DeepSeek-V3-671B, which exhibits a strong tendency to assign a score of 5 without bias. However, the introduction of a reference answer scored at 5 mitigates this phenomenon.
In summary, the above analysis reveals that the score of the reference answer introduced in the prompt has a significant impact on the judge model’s scoring behavior. In particular, when reference answers with a score of 5 are introduced, the models achieve the highest scoring accuracy without disproportionately influencing their scoring preferences. Therefore, it can be concluded that when incorporating reference answers into the scoring-based judge prompt, a full-marked reference answer is the best choice. Interestingly, this conceptually aligns with the principles of traditional statistical metrics, such as BLEU and ROUGE, which emphasize the critical role of golden answers in evaluation.
Limitations and Future Work
Although our work offers notable conclusions and insights, some limitations still remain for future work.
Mitigating scoring bias needs further exploration. Although there are no dedicated and fancy methods for addressing scoring bias, intuitive and straightforward approaches, such as scoring multiple times followed by majority voting or averaging the scores, require empirical validation for their effectiveness in mitigating scoring bias.
Additionally, although our work demonstrates the existence of scoring bias in LLM-as-a-Judge, the underlying causes of scoring bias remain to be validated. Approaches such as training data analysis and information flow observation (Yuan et al. 2025) may help identify the reason scoring bias exists within or outside the model.
Conclusion
In this work, we propose scoring bias in LLM-as-a-Judge, defined as scoring judgments shift when judge models are perturbed. We design an evaluation framework, including three types of scoring bias: score rubrics order bias, score IDs bias, and reference answer score bias. We select four LLM-as-a-Judge benchmarks and expand them as our evaluation datasets through data synthesis. Our evaluation metrics can quantify the impact of scoring bias on the scoring accuracy and scoring tendencies of the judge models. Our experimental results clearly demonstrate that scoring bias can affect the robustness and reliability of existing scoring-based judge models. Judge models with different capabilities, sizes, and architectures exhibit different sensitivities to various scoring biases. Our further exploratory experiments and discussions reveal several intriguing phenomena. In summary, our work provides detailed evidence and valuable recommendations for evaluating and improving the scoring reliability of LLM-as-a-Judge.